Fermi National Laboratory
 

Accelerator Update for February 9, 2007 - February 12, 2007

 

Friday February 9

The day shift began with Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (303.3E10) and the Antiproton Source (Pbar) stack (17.3E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the colliding beams.

Operations put the TeV into shot setup at 8:10 AM.

At 8:23 AM, Operations reported that the 750keV line buncher had tripped off. Operators investigated.

At 8:44 AM, the NuMI beam permit tripped off due to a bad beam position. It reset okay.

At 9:03 AM, the P1 transfer line tripped off and took down NuMI with it. Everything reset okay.

At 9:35 AM, Operations reported that NuMI had been tripping off due to bad beam positions. Operations reduced the beam intensity to NuMI and a MI expert began tuning.

RF experts began preparing a Booster RF station (BRF6) at 10:40 AM to replace its modulator.

Operations established store 5220 at 10:50 AM with an initial luminosity of 252.23E30.

At 11:10 AM, Operations reported that something had happened to the MI beam positions at 9:12 AM that caused the NuMI autotune program to make a big change. When the MI beam positions returned to normal a couple of pulses later, the autotune had already gone beyond its operational limits and stopped tuning. (This may have been due to a kicker problem.) Everything reset okay.

Switchyard (SY) experts began a study period at 12:01 PM.

Operations reported at 1:37 PM that the Booster beam was turned off to tie BRF6 back into the system.

Operations turned off the SY beam at 1:38 PM and prepared the Meson enclosures M03, 4, & 5 for access.

Booster beam resumed at 2:11 PM with BRF6 back online. Pbar and NuMI resumed taking beam.

MI began a NuMI 11- batch slipstacking study at 3:27 PM. The study ended at 4:27 PM. These studies yielded a new MI intensity record of 3.5E13 protons at 120 GeV.

Operations reported at 4:14 PM that the Recycler had stashed antiprotons one time during the day shift.

At 5:23 PM, BRF6 tripped off for the first of 13 times in 4.5 hours. Operations reset the station each time. After the last trip, an RF expert asked Operations to make a voltage adjustment.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Saturday February 10

The midnight shift began with Operations monitoring store 5220 (65.63E30) and the stash (173.1E10), with Pbar stacking (39.1E10), and with NuMI taking beam.

At 1:03 AM, BRF6 tripped off. It reset.

At 1:35 AM, the Booster chopper tripped off. Operators investigated and discovered it was sparking. Operations called in experts.

Booster experts completed their repairs on the chopper at 4 AM. They reported that a cable had shorted and was replaced.

At 5:33 AM, Operations reported that the Electron Cooling system node had gone down. Operators talked with experts and then accessed MI-31 to give it a local reboot. This didn’t resolve the problem. Operations called in experts.

At 6:33 AM, the Recycler tune began to drift. Operations contacted experts and with their help recovered the tune. Unfortunately, about 6E10 of the stash was lost.

Recycler experts brought Electron Cooling system back on line at 8:43 AM.

A TeV expert began an end-of-store study at 12:02 PM. A little later another TeV expert began a crystal collimator study.

Operations terminated store 5220 at 1:56 PM.

Operations put the TeV into shot setup at 2:10 PM.

At 4:16 PM, Electron Cooling system tripped off. It reset and resumed normal operations at 4:29 PM.

Operations established store 5221 at 5:43 PM with an initial luminosity of 149.99E30.

Recycler experts accessed MI at 6:21 PM. They were out 20 minutes later. They had replaced a regulator for their cooling system (R:H4A21).

Operations reported that the Recycler had completed one stash around 8 PM.

At 10:48 PM, at TeV RF station (TRF1) tripped off. Experts will check the station when this store ends.

Sunday February 11

The midnight shift began with Operations monitoring store 5221 and the stash, with Pbar stacking, and with NuMI taking beam.

At 3:36 AM, LRF2 tripped off. It reset.

At 4:12 AM, the NuMI beam permit tripped off due to a bad beam position. It reset.

At 4:26 AM, a Pbar Debuncher RF station (DRF1) tripped off for the first of five times in two hours. It reset each time.

Operations reported at 8:03 AM that the Recycler had stashed five times since 8 PM on Saturday.

At 8:20 AM, Operations reported that Electron Cooling system node wasn’t responding again. An expert had operations reboot the system twice, but it didn’t help.

At 9:39 AM, LRF5 tripped off. It reset.

At 10:20 AM, DRF1 tripped off. Operators could not reset the station. They contacted experts.

A Recycler expert had Electron Cooling system node working by 10:56 AM.

Operations terminated store 5221 at noon.

At 12:02 PM, LRF5 tripped off. It reset.

An RF expert reported at 2:04 PM that he had changed out TRF1’s modulator, but then found a LCW leak on the Power Amplifier.

At 2:15 PM, Operations reported that a cryo system expert was coming in to investigate a problem with the TeV sector A0 cryo system.

At 3:09 PM, LRF5 tripped off. It reset.

At 4:23 PM, Electron Cooling system tripped off again. The Pelletron went through a full discharge. Operators called an expert who advised Operators on how to recover the system. Normal cooling resumed about an hour later.

TeV experts completed their repairs on TRF1 at 4:41 PM.

Operations put the TeV through a dry squeeze at 4:49 PM.

At 5:07 PM, Operations reported that there was still a problem with DRF1. Experts came in to investigate.

At 5:31 PM, the Pelletron discharged again. Operations called in an expert.

Operations prepared D0 for an access at 6:01 PM.

At 6:34 PM, a Booster vacuum pump tripped off and caused BRF15 & 16 to trip. Everything reset okay.

At 6:47 PM, BRF6 tripped off. It reset.

At 7:16 PM, D0 reported that around 4:30 PM they were hit with a power glitch. They had a VESDA alarm and a number of devices had tripped off. The Duty Electrician investigated. Operations didn’t see anything resembling a power glitch.

Operations reported at 8:53 PM that they had turned on DRF1.

At 11:44 PM, Operations called in a Pbar expert to help with emittance problems.

Monday February 12

The midnight shift began with Operations monitoring the stash (263.9E10), with Pbar stacking (176.5E10), and with NuMI taking beam.

Operations reported the following status at 12:48 AM:

  • The Pbar stack was at 200mA and was having stability problems
  • The Recycler stash was at 262E10 and cooling to prepare for a stash
  • The TeV was at 150 GeV and waiting for D0 to come out of access
  • D0 was experiencing controls problems

At 3:49 AM, BRF6 tripped off. It reset.

D0 came out of access at 5:31 AM. They reported that everything looked okay.

Operations put the TeV through a dry squeeze at 5:58 AM.

Operations put the TeV into shot setup at 6:31 AM.

Operations reported at 8 AM that the Recycler had stashed once during the midnight shift.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar experts to access AP1 and work on their target this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/12/2007    email Fermilab